OpenStudy (kalli):

can someone explain to me how this is correct? This is an answer my teacher gave me Find the exact value for Tan(17pi/12) Answer ---> ( sqrt3 +1)/(sqrt3 - 1)

OpenStudy (alyssa_xo):

you can expess tan as sin/cos

OpenStudy (anonymous):

this is a time to use a scientific calculator

OpenStudy (alyssa_xo):

\(\sin (\dfrac{17 \pi}{12})=-\dfrac{\sqrt3+1}{2\sqrt2}\) find cosine, simplify the fraction. I assume the denominators cancel out

OpenStudy (kalli):

o, that was so simple haha i wasnt sure. Thanks @alyssa_xo

OpenStudy (alyssa_xo):

anytime babygurl